OpenStudy (anonymous):

The freshman class has 160 students. Each freshman must write three reports from a list of 12 novels that the teacher has assigned them to read. How many different combinations of novels can a freshman student choose to write reports on?

OpenStudy (anonymous):

you are being asked for \[\binom{12}{3}\] sometimes written as \[_{12}C_3\] or even \[^{12}C_3\] and computed via \[\binom{12}{3}=\frac{\overbrace{12\times 11\times 10}^{\text{ 3 terms}}}{3\times 2}\]

OpenStudy (anonymous):

cancel first, multiply last, since the answer is obviously a whole number